• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:26:32 on December 6,2007

    Entry number 218445

    keyword=no beam

    We have not had any production beam yet since coming out of controlled access ~4:30 pm. There were only occational beam for tuning, and checking. No beam now since MCC is trouble shooting Hall C beam problem. When trying to delieve high current to Hall C, it caused valve(s) to close (in South LINAC?). They are discussing with optics-on-call.